OpenStudy (anonymous):

What is the recursive formula for 1, 4, 13, 40, 121, …?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

will someone please help me with recursive formula i just want to know how it works

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@Zale101

OpenStudy (anonymous):

@dan815

OpenStudy (zarkon):

4=3*1+1

OpenStudy (anonymous):

what?

OpenStudy (zarkon):

13=3*4+1

OpenStudy (anonymous):

what do you mean im coonfused

OpenStudy (anonymous):

\[ a_2 = 3a_1+1,\quad a_3 = 3a_2+1,\quad a_4 = 3a_3+1 \]

OpenStudy (anonymous):

what are you trying to tell me??

OpenStudy (zarkon):

you might want to review what a recursive formula is

OpenStudy (anonymous):

it makes no sense to me

OpenStudy (anonymous):

wio will try to help me understand this

OpenStudy (anonymous):

If each term of a sequence changes by a common difference, d, what would a general rule for this type of sequence be?

OpenStudy (anonymous):

i have no idea what that mean at all

OpenStudy (anonymous):

\[\Large a_2 = 3a_1 + 1 \implies a_2= 3a_{2-1} +1 \]